OpenStudy (anonymous):

Use the quadratic formula to solve 4y^2 + 8y + 7 = 4

OpenStudy (solomonzelman):

\[4y^2+8y+7=4\]\[4y^2+8y+7\color{red} { -4 } =4\color{red} { -4 }\]\[4y^2+8y+3=0\]look for the a b c and plug them into the quadratic formula.\[\color{blue} { 4 }y^2+\color{blue} { 8 }y+\color{blue} { 3 }=0\]\[\color{green} { a=4 ~~~~~~~ }\color{green} { b=8 ~~~~~~~ }\color{green} { c=3 ~~~~~~~ }\]

OpenStudy (solomonzelman):

\[\huge\color{red}{ \frac{-\huge\color{blue}{ 8 } ±\sqrt{\huge\color{blue}{ 8 } ^2-(4 \times \huge\color{blue}{ 4 } \times \huge\color{blue}{ 7 } )}}{2 \times \huge\color{blue}{ 4} }}\]

OpenStudy (solomonzelman):

\[\huge\color{goldenrod}{ \frac{-8±\sqrt{-48}}{8} }\]

OpenStudy (solomonzelman):

if you haven't learned about imaginary number (I am talking about "i") then simply say no solution. Now, if you have lets go on... \[\huge\color{goldenrod}{ \frac{-8±4i \sqrt{3}}{8} }\] and divide top and bottom by 4, \[\huge\color{cornflowerblue}{ \frac{-2±i \sqrt{3}}{2} }\]